# Semantic Scene Completion Based 3D Traversability Estimation for Off-Road Terrains

### v1

@@ III-A1 Feature Extraction
To achieve more accurate occupancy estimations, multimodal information is utilized to extract environmental features. The model employs a ResNet-50 backbone to extract image features $F_{t}^{2D}\in\mathbb{R}^{l\times b\times d}$ from the monocular camera image $I_{t}$, where $l\times b$ represents the spatial resolution and $d$ denotes the feature dimension. Concurrently, the sparse point clouds $L_{t}$ from the LiDAR input are voxelized to generate sparse voxel occupancy supervision $S_{\text{spa}}\in\{0,1\}^{H\times W\times Z}$. Subsequently, a query upscale network is utilized to produce lower spatial resolution occupancy supervision $S_{\text{den}}\in\{0,1\}^{H/2\times W/2\times Z/2}$, thereby enhancing the robustness of the queries.
Mask Token. The mask token module proposed by Y. Li [56] is incorporated to represent empty voxel spaces within the environment. When combined with voxel occupancy supervision, this approach forms a more comprehensive 3D voxel feature, facilitating improved occupancy predictions.
@@ III-A2 Feature Fusion
In the model, deformable attention (DA) mechanisms [57] are adopted to facilitate the interaction between local regions of interest in the 3D and 2D feature spaces. Deformable attention dynamically samples $N_{s}$ points around a reference point, and the attention weights are adaptively learned based on the query vector, while the spatial offsets for sampled points are predicted relative to the reference point, enabling flexible and efficient feature aggregation. The attention results are computed based on these samples using the following equation:
@@
where, $q$ is the query vector, $p$ denotes the reference point associated with $q$, $F$ represents the input feature map, $W_{s}\in\mathbb{R}^{d\times d}$ are learnable weight matrices for value generation at each sampled point, $A_{s}\in[0,1]$ are attention weights for each sampled point, dynamically learned based on the query, $\delta p_{s}\in\mathbb{R}^{2}$ are the predicted offsets for the sampled points relative to the reference point $p$, and $F(p+\delta p_{s})$ denotes the feature vector at the location $p+\delta p_{s}$, extracted via bilinear interpolation from the input features.
Deformable Cross-Attention (DCA). Multiple DCA layers are employed to enhance the interaction between the occupancy supervision queries $S_{\text{den}}$ and the image features $F_{t}^{2D}$, thereby facilitating the integration of multimodal information. The 3D grids are projected onto the 2D image feature space $F_{t}^{2D}$ using projection matrices. The resulting 2D points serve as reference points for the queries $S_{\text{den}}$, from which surrounding features are sampled. A weighted sum of the sampled features is computed to produce the output of the DCA layers. Consequently, refined queries $\hat{S}_{\text{den}}$ are obtained, fusing environmental semantic features with geometric features.
@@
Deformable Self-Attention (DSA). The refined queries $\hat{S}_{\text{den}}$ are concatenated with mask tokens to form the initial voxel features $F^{3D}\in\mathbb{R}^{H/2\times W/2\times Z/2\times d}$. Subsequently, these features are processed through multiple DSA layers to obtain the enhanced voxel features $\hat{F}^{3D}\in\mathbb{R}^{H/2\times W/2\times Z/2\times d}$. The DSA layers enable the model to learn spatial 3D occupancy features more efficiently and accurately. In the equations, the query $q_{p}$ can be either a mask token or a refined query located at position $p=(x,y,z)$.
@@ III-A3 Output
The enhanced voxel features $\hat{F}^{3D}$ are input to fully connected (FC) layers and passed through a softmax function to produce dense semantic occupancy predictions $Y_{t}\in\{c_{0},c_{1},c_{2},\cdots,c_{M}\}^{H\times W\times Z}$, where each voxel is classified as either empty $\{c_{0}\}$ or occupied by a specific traversability class within $\{c_{1},c_{2}\cdots,c_{M}\}$. Mapping these predictions back to the 3D grid yields a detailed occupancy map that highlights traversable areas and provides semantic labels essential for autonomous off-road navigation.
@@ III-A4 Training Loss
ORDformer is trained end-to-end from scratch by minimizing three primary loss components: the semantic scale loss $\mathcal{L}_{\text{scal}}^{\text{sem}}$, the geometric scale loss $\mathcal{L}_{\text{scal}}^{\text{geo}}$, and the weighted cross-entropy loss $\mathcal{L}_{\text{ce}}$. The total training loss $\mathcal{L}_{\text{total}}$ is defined as the sum of these individual losses:
@@
In the context of semantic scene completion, model performance is evaluated using several metrics, including Precision $P_{c}$, Recall $R_{c}$, and Specificity $S_{c}$. These metrics assess the ability of the model to correctly classify occupied and non-occupied voxels. The metrics are formulated as follows:
@@
III-B 3D Traversable Occupancy Annotations
The RELLIS-OCC dataset with 3D traversable occupancy annotations is generated based on the RELLIS-3D dataset [18], comprising five sequences and a total of 11,201 frames. In this context, sequences numbered 00000 to 00003 are designated as the training set, while sequence 00004 is utilized as the validation set.
Figure 3: Dense occupancy ground truth generation.
III-B1 Dense Semantic Occupancy Annotations
In the experiments, networks trained with sparse LiDAR points fail to predict sufficiently dense occupancy, highlighting the need for dense occupancy labels. Dense forward-facing labels are generated from the semantically annotated point clouds of the RELLIS-3D dataset, as shown in Fig. 3. The method transforms multi-frame LiDAR point sequences into a unified coordinate system and voxelizes the dense points into grids, including dynamic objects. This process also captures surface occupancy information since LiDAR scans surface points.
Figure 4: Voxel Semantic Label Distribution: Grass, trees, and bushes constitute the major classes among the nine valid semantic classes. Each category’s color corresponds to its semantic label.
Due to the complex and occluded nature of off-road obstacles, occupancy ground truth within a range of $[38.4\,\mathrm{m},51.2\,\mathrm{m},8\,\mathrm{m}]$ is selected. The prediction range is set to $[0\,\mathrm{m},38.4\,\mathrm{m}]$ along the X-axis, $[-25.6\,\mathrm{m},25.6\,\mathrm{m}]$ along the Y-axis, and $[-2\,\mathrm{m},6\,\mathrm{m}]$ along the Z-axis. The ground truth is voxelized into a $192\times 256\times 40$ grid with a voxel size of $0.2\,\mathrm{m}$ and data is filtered to retain voxels within the camera’s field of view. A remapping of the original categories is then performed to assess their impact on traversability. Voxel data statistics for the RELLIS-OCC dataset, comprising 10 semantic category labels (including void), are presented in Fig. 4.
@@ III-B2 Geometric Feature Parameter Calculation
Step Height $h$: The maximum elevation difference between the central voxel $G$ and other voxels $G_{j}$ within its neighborhood $\mathrm{\Omega}$ is calculated to assess the variation in terrain height around the central voxel.
@@
where, $G^{z}$ and $G_{j}^{z}$ represent the elevations of the central voxel $G$ and the neighboring voxel $G_{j}$, respectively.
@@
where, the value of $\mathbf{n}$ is obtained by calculating the covariance matrix using Principal Component Analysis (PCA).
@@
where, $a_{0}$ and $a_{1}$ are the coefficients representing the slope in the $x$ and $y$ directions, respectively, and $c$ is the intercept, $z_{\text{actual}}$ represents the actual elevation of the current voxel.
@@ III-B3 Vehicle Obstacle Crossing Condition Analysis
In off-road conditions, the traversable area of a vehicle is strongly coupled with its geometric passability, as the difficulty of traversing the ground relates to various road geometric features. An analysis of the failure conditions for vehicle obstacle crossing, based on the principles of vehicle ground mechanics, results in the creation of the Step Mask.
Step Mask: The failure conditions for vehicle obstacle crossing include body suspension, front-end collision, vertical obstacles, ditches, and lateral tipping. Since the dataset does not encompass particularly challenging scenarios, the analysis is focused on four specific cases: overcoming vertical obstacles, crossing ditches, crossing overhang obstacles and traversing longitudinal slopes[58].
@@
The performance of wheeled vehicles in overcoming step-like obstacles and vertical protrusions is primarily evaluated based on the maximum height of vertical obstacles they can surmount. The dimensionless expression for the maximum obstacle height that the front wheels of a 4×4 wheeled vehicle can overcome is:
$\frac{h}{r}=\frac{1-\mu\frac{r}{l}+\eta^{2}-\eta\sqrt{1-2\mu\frac{r}{l}+\eta^{2}}}{\left(1+\mu\frac{r}{l}\right)^{2}+\eta^{2}}.$ (14)
$\eta=\frac{1-\mu\frac{r}{l}-(1+\mu^{2})\frac{a}{l}}{\mu}.$ (15)
where, $h$ represents the height of the step obstacle, $r$ represents the wheel radius, $\mu$ represents the coefficient of friction, $l$ represents the wheelbase of the vehicle, $a$ represents the horizontal distance from the front axle to the center of gravity of the vehicle.
@@
The width of the trenches that a wheeled vehicle can overcome is evaluated using the ratio of the trench width $l_{d}$ to the wheel diameter $D$, expressed as $\frac{l_{d}}{D}$. The ratio of the height of vertical obstacles that can be overcome to the wheel diameter, expressed as $\frac{h}{D}$, can be converted to $\frac{l_{d}}{D}$, as shown in Equation16. Given a specific obstacle height $h$, the corresponding trench width ratio $\frac{l_{d}}{D}$ can be determined.
$\frac{l_{d}}{D}=2\sqrt{\frac{h}{D}-\left(\frac{h}{D}\right)^{2}}.$ (16)
@@
For a vehicle to successfully traverse overhang obstacles, the height $h_{\text{obj}}$ of the lowest point of the obstacle above the ground must exceed the sum of $h_{\text{pc}}$, which is the height of the obstacle point cloud relative to the LiDAR sensor, and $h_{\text{Lid}}$, the height of the LiDAR sensor above the ground. When the vehicle encounters an overhang obstacle, the following conditions must be met:
$h_{\text{obj}}>h_{\text{pc}}+h_{\text{Lid}}.$ (17)
@@
For traversing longitudinal slopes, the ground platform must satisfy the following condition:
$\theta_{\text{max}}>\alpha,$ (18)
where $\theta_{\text{max}}$ represents the maximum climbing angle that the vehicle can handle, and $\alpha$ represents the current slope angle of the terrain.
III-B4 Traversability Cost Occupancy Annotations
@@
Rules are established to create an initial mapping from voxel semantic categories to traversability cost, with adjustments made based on the previously computed Step Mask. The Step Mask is determined by evaluating the geometric passability of the vehicle. Voxels not meeting the geometric passability criteria are assigned a lethal cost. Finally, four traversability cost labels are derived: lethal, medium-cost, low-cost, and free, ranging from difficult to easy traversability. Fig. 5 illustrates the statistical distribution of the traversability cost labels.
Figure 5: Voxel traversability cost label distribution: Each category’s color corresponds to its traversability cost label.

### v2

@@ III-A1 Feature Extraction
To achieve more accurate terrain modeling, multimodal information is utilized to extract environmental features. The model employs a ResNet-50 backbone to extract image features $F_{t}^{2D}\in\mathbb{R}^{l\times b\times d}$ from the monocular camera image $I_{t}$, where $l\times b$ represents the spatial resolution and $d$ denotes the feature dimension. Concurrently, the sparse point clouds $L_{t}$ from the LiDAR input are voxelized to generate sparse voxel supervision $S_{\text{spa}}\in\{0,1\}^{H\times W\times Z}$. Subsequently, a query upscale network is utilized to produce lower spatial resolution voxel supervision $S_{\text{den}}\in\{0,1\}^{H/2\times W/2\times Z/2}$, thereby enhancing the robustness of the queries.
Mask Token. The mask token module proposed by Y. Li [50] is incorporated to represent empty voxel spaces within the environment. When combined with voxel supervision, this approach will form a more comprehensive 3D voxel feature, facilitating improved occupancy predictions.
@@ III-A2 Feature Fusion
In the model, deformable attention (DA) mechanisms [51] are adopted to propel the interaction between local regions of interest in the 3D and 2D feature spaces. Deformable attention dynamically samples $N_{s}$ points around a reference point, and the attention weights are adaptively learned based on the query vector, while the spatial offsets for sampled points are predicted relative to the reference point, enabling flexible and efficient feature aggregation. The attention results are computed based on these samples using the following equation:
@@
where $q$ is the query vector, $p$ denotes the reference point associated with $q$, $F$ represents the input feature map, $W_{s}\in\mathbb{R}^{d\times d}$ are learnable weight matrices for value generation at each sampled point, $A_{s}\in[0,1]$ are attention weights for each sampled point, dynamically learned based on the query, $\delta p_{s}\in\mathbb{R}^{2}$ are the predicted offsets for the sampled points relative to the reference point $p$, and $F(p+\delta p_{s})$ denotes the feature vector at the location $p+\delta p_{s}$, extracted via bilinear interpolation from the input features.
Deformable Cross-Attention (DCA). Multiple DCA layers are employed to enhance the interaction between the occupancy supervision queries $S_{\text{den}}$ and the image features $F_{t}^{2D}$, thereby reinforcing the integration of multimodal information. The 3D grids are projected onto the 2D image feature space $F_{t}^{2D}$ using projection matrices. The resulting 2D points serve as reference points for the queries $S_{\text{den}}$, from which surrounding features are sampled. A weighted sum of the sampled features is computed to produce the output of the DCA layers. Consequently, refined queries $\hat{S}_{\text{den}}$ are obtained, fusing environmental semantic features with geometric features.
@@
Deformable Self-Attention (DSA). The refined queries $\hat{S}_{\text{den}}$ are concatenated with mask tokens to form the initial voxel features $F^{3D}\in\mathbb{R}^{H/2\times W/2\times Z/2\times d}$. Subsequently, these features are processed through multiple DSA layers to obtain the enhanced voxel features $\hat{F}^{3D}\in\mathbb{R}^{H/2\times W/2\times Z/2\times d}$. The DSA layers enable the model to learn spatial 3D occupancy features more efficiently and accurately. In the equations, the query $q_{p}$ can either be a mask token or a refined query located at position $p=(x,y,z)$.
@@ III-A3 Output
The enhanced voxel features $\hat{F}^{3D}$ are input to fully connected (FC) layers and passed through a softmax function to produce dense traversability predictions $Y_{t}\in\{c_{0},c_{1},c_{2},\cdots,c_{M}\}^{H\times W\times Z}$, where each voxel is classified as either empty $\{c_{0}\}$ or occupied by a specific traversability class within $\{c_{1},c_{2}\cdots,c_{M}\}$. Mapping these predictions back to the 3D grid yields a detailed occupancy map that highlights traversable areas and provides traversability labels essential for autonomous off-road navigation.
@@ III-A4 Training Loss
3DTTNet is trained end-to-end from scratch by minimizing three primary loss components: the semantic scale loss $\mathcal{L}_{\text{scal}}^{\text{sem}}$, the geometric scale loss $\mathcal{L}_{\text{scal}}^{\text{geo}}$, and the weighted cross-entropy loss $\mathcal{L}_{\text{ce}}$. The total training loss $\mathcal{L}_{\text{total}}$ is defined as the sum of these individual losses:
@@
In the context of semantic scene completion, model performance is evaluated via several metrics, including Precision $P_{c}$, Recall $R_{c}$, and Specificity $S_{c}$. These metrics assess the ability of the model to correctly classify occupied and non-occupied voxels. The metrics are formulated as follows:
@@
III-B 3D Traversable Annotations
We propose a labeling pipeline that integrates semantic accessibility analysis with geometric passability considerations, resulting in the generation of the RELLIS-OCC dataset. This dataset, derived from the semantically annotated point clouds of the RELLIS-3D dataset [15], comprises five sequences and a total of 11,201 frames. Sequences 00000 to 00003 are designated as the training set, while sequence 00004 is used for validation.
III-B1 Dense Semantic Annotations
In the experiments, networks trained with sparse LiDAR points fail to predict sufficiently dense occupancy, highlighting the need for dense occupancy labels. Dense forward-facing labels are generated from the semantically annotated point clouds of the RELLIS-3D dataset. We transform multi-frame LiDAR scans into a unified coordinate system and voxelize the dense points into grids, including dynamic objects. This process captures surface information since LiDAR measures points on object surfaces.
Figure 3: Voxel Label Distribution: (a) the distribution of semantic annotations, and (b) the distribution of traversability cost annotations. Each category’s color corresponds to its label.
Due to the complex and occluded nature of off-road obstacles, ground truth within a range of $38.4\,\mathrm{m},51.2\,\mathrm{m},8\,\mathrm{m}]$ is selected. The prediction range is set to $0\,\mathrm{m},38.4\,\mathrm{m}]$ along the X-axis, $25.6\,\mathrm{m},25.6\,\mathrm{m}]$ along the Y-axis, and $2\,\mathrm{m},6\,\mathrm{m}]$ along the Z-axis. The ground truth is voxelized into a $192\times 256\times 40$ grid with a voxel size of $0.2\,\mathrm{m}$ and data is filtered to retain voxels within the camera’s field of view. A remapping of the original categories is then performed to assess their impact on traversability. Voxel data statistics of the RELLIS-OCC dataset, comprising 11 semantic category labels (including void), are presented in Fig. 3(a).
@@ III-B2 Geometric Feature Parameter Calculation
Step Height $h$: The maximum elevation difference between the central voxel $G$ and other voxels $G_{j}$ within its neighborhood $\mathrm{\Omega}$ is calculated, in order to assess the variation in terrain height around the central voxel.
@@
where $G^{z}$ and $G_{j}^{z}$ represent the elevations of the central voxel $G$ and the neighboring voxel $G_{j}$, respectively.
@@
where the value of $\mathbf{n}$ is obtained by calculating the covariance matrix using Principal Component Analysis (PCA).
@@
where $a_{0}$ and $a_{1}$ are the coefficients representing the slope in the $x$ and $y$ directions, respectively, $c$ is the intercept, and $z_{\text{actual}}$ represents the actual elevation of the current voxel.
@@ III-B3 Vehicle Obstacle Crossing Condition Analysis
In off-road conditions, the traversable area of a vehicle is strongly coupled with its geometric passability, as the difficulty of traversing the ground relates to various road geometric features. By analyzing the failure conditions during vehicle obstacle negotiation, we introduce the Geometric Passability Assessment Mask (GPAM).
GPAM: The failure conditions for vehicle obstacle crossing are composed of body suspension, front-end collision, vertical obstacles, ditches, and lateral tipping. Since the dataset does not encompass particularly challenging scenarios, the analysis is focused on four specific cases: overcoming vertical obstacles, crossing ditches, crossing overhang obstacles and traversing longitudinal slopes [52].
@@
The performance of wheeled vehicles in overcoming step-like obstacles and vertical protrusions is primarily evaluated based on the maximum height of vertical obstacles they can surmount. The dimensionless expression for the maximum obstacle height $h_{\text{max}}$ that the front wheels of a 4×4 wheeled vehicle can overcome is:
$\frac{h_{\text{max}}}{r}=\frac{1-\mu\frac{r}{l}+\eta^{2}-\eta\sqrt{1-2\mu\frac{r}{l}+\eta^{2}}}{\left(1+\mu\frac{r}{l}\right)^{2}+\eta^{2}},$ (14)
$\eta=\frac{1-\mu\frac{r}{l}-(1+\mu^{2})\frac{a}{l}}{\mu},$ (15)
where $r$ denotes the wheel radius, $\mu$ signifies the coefficient of friction, $l$ refers to the wheelbase of the vehicle, and $a$ corresponds to the horizontal distance from the front axle to the center of gravity of the vehicle. Thus, the condition for a vehicle to overcome a vertical obstacle is expressed as follows:
$\Psi_{\text{ver}}(G_{j})=\left\llbracket h\leq h_{\text{max}}\right\rrbracket.$ (16)
@@
The width of the trenches that a wheeled vehicle can overcome is evaluated using the ratio of the trench width $l_{d}$ to the wheel diameter $D$, expressed as $\frac{l_{d}}{D}$. The ratio of the height of vertical obstacles $h_{d}$ that can be overcome to the wheel diameter, expressed as $\frac{h_{d}}{D}$, can be converted to $\frac{l_{d}}{D}$, as shown in Equation17. Given a specific obstacle height, the corresponding trench width ratio $\frac{l_{d}}{D}$ can be determined.
$\frac{l_{d}}{D}=2\sqrt{\frac{h_{d}}{D}-\left(\frac{h_{d}}{D}\right)^{2}}.$ (17)
Therefore, the condition for a vehicle to overcome a trench, where $l_{d{\text{max}}}$ represents the maximum trench width the vehicle can cross, calculated when $h_{d}$ equals $h_{\text{max}}$ using Equation17, is expressed as:
$\Psi_{\text{tre}}(G_{j})=\left\llbracket l_{d}\leq l_{d_{\max}}\right\rrbracket.$ (18)
@@
To ensure a vehicle successfully traversing overhang obstacles, the height $h_{\text{obj}}$ of the lowest point of the obstacle above the ground must exceed the sum of $h_{\text{pc}}$ (which is the height of the obstacle point cloud relative to the LiDAR sensor), and $h_{\text{Lid}}$ (which is the height of the LiDAR sensor above the ground). Once the vehicle encounters an overhang obstacle, the following conditions must be met:
$\Psi_{\text{ove}}(G_{j})=\left\llbracket h_{\text{obj}}>h_{\text{pc}}+h_{\text{Lid}}\right\rrbracket.$ (19)
@@
In off-road conditions, the maximum slope angle $\emptyset_{m}$ that a vehicle can traverse is governed by the terrain’s friction-related angle $\emptyset_{\mu}$ and the vehicle’s approach angle $\emptyset_{a}$. The friction-related angle $\emptyset_{\mu}$ can be calculated using:
$\emptyset_{\mu}=\sin^{-1}\left(\frac{F_{\text{max}}}{mg\sqrt{\mu_{s}^{2}+1}}\right)-\tan^{-1}(\mu_{s}),$ (20)
where $F_{\text{max}}$ is the vehicle’s maximum propulsion force, $m$ is the vehicle mass, $g$ is gravitational acceleration, and $\mu_{s}$ is the terrain’s unevenness parameter. This formula accounts for both surface roughness and traction limitations [53]. The ultimate traversable slope angle $\emptyset_{m}$ is then determined by the limiting factor between terrain friction and vehicle geometry:
$\emptyset_{m}=\min(\emptyset_{\mu},\emptyset_{a}),$ (21)
Therefore, the condition for a vehicle to overcome a slope is expressed as:
$\Psi_{\text{slo}}(G_{j})=\left\llbracket s\leq\emptyset_{m}\right\rrbracket.$ (22)
Above, the expression of GPAM is:
$\text{GPAM}(G_{j})=\left\llbracket\Psi_{\text{ver}}(G_{j})\land\Psi_{\text{tre}}(G_{j})\land\Psi_{\text{ove}}(G_{j})\land\Psi_{\text{slo}}(G_{j})\right\rrbracket.$ (23)
III-B4 Traversability Cost Annotations
@@
To establish an initial mapping from voxel semantic categories to traversability cost, a set of rules is defined, incorporating adjustments based on the previously computed GPAM, which evaluates the vehicle’s geometric passability. Voxels failing to meet the required passability criteria are assigned a lethal cost. This results in four distinct traversability cost labels: lethal, medium-cost(M-cost), low-cost(L-cost), and free, which represent a spectrum of traversability difficulty, from challenging to easy. The statistical distribution of these labels is shown in Fig. 3(b).
The traversability cost labeling function $F(G_{j})$ is formalized as follows:
$F(G_{j})=\begin{cases}\text{Lethal},&\text{if }S(G_{j})\in S_{\text{lethal}}\ \lor\ \text{GPAM}(G_{j})=0\\ \text{M-cost},&\text{if }S(G_{j})\in S_{\text{traversable}}\ \land\\ &\phi_{\text{geo}}(G_{j})\geq\tau_{\text{medium}}(V)\\ \text{L-cost},&\text{if }S(G_{j})\in S_{\text{traversable}}\ \land\\ &\tau_{\text{low}}(V)\leq\phi_{\text{geo}}(G_{j})<\tau_{\text{medium}}(V)\\ \text{Free},&\text{if }S(G_{j})\in S_{\text{free}}\ \land\ \phi_{\text{geo}}(G_{j})<\tau_{\text{low}}(V)\end{cases}$ (24)
Semantic Mapping
The semantic categories are divided into three types: lethal obstacles that the vehicle cannot traverse ($S_{\text{lethal}}=\{\text{person},\text{object},\text{barrier},\text{tree}\}$), traversable terrains that require verification ($S_{\text{traversable}}=\{\text{water},\text{soil},\text{bush},\text{grass}\}$), and free surfaces considered inherently traversable ($S_{\text{free}}=\{\text{hard-surface},\text{mud},\text{rubble}\}$).
Geometric Risk Evaluation Function
The geometric risk evaluation function $\phi_{\text{geo}}(G_{j})$ quantifies the passability of each voxel based on several terrain characteristics:
$\phi_{\text{geo}}(G_{j})=\frac{h}{h_{\max}(V)}+\frac{s}{s_{\max}(V)}+\frac{u}{u_{\text{critic}}(V)},$ (25)
where $h_{\max}(V)$ is the maximum vertical obstacle height that the vehicle can overcome (Equation 14), $s_{\max}(V)$ represents the maximum slope angle the vehicle can traverse (Equation 21), and $u_{\text{critic}}(V)=\log\left(\min\left(z_{\text{susp}},{z_{\text{clearance}}}\right)\right)$ is the critical unevenness, which depends on the suspension displacement $z_{\text{susp}}$ and the minimum ground clearance $z_{\text{clearance}}$ of the vehicle chassis.
Threshold Calibration
Thresholds for classifying terrain into medium-cost and low-cost categories are dynamically derived based on vehicle parameters:
$\tau_{\text{medium}}(V)=\eta_{\text{medium}}\times\phi_{\text{geo, max}}(V),$ (26)
$\tau_{\text{low}}(V)=\eta_{\text{low}}\times\phi_{\text{geo, max}}(V),$ (27)
where $\eta_{\text{medium}}$ and $\eta_{\text{low}}$ are scaling factors that adjust the thresholds for medium and low-cost classifications, respectively. $\phi_{\text{geo, max}}(V)$ represents the theoretical maximum geometric risk when $h=h_{\max},s=s_{\max},u=u_{\text{critic}}$.
In conclusion, the pipeline for generating traversability cost annotations is outlined in Algorithm 1.
Algorithm 1 Labeling Pipeline
0: Point clouds $g$, Voxel grid $G$, Vehicle parameters $V$
0: Traversability labels $F(G_{j})$
1: Initialize semantic mapping:
[…]
5:for each voxel $G_{j}$ do
6:  if $S(G_{j})\in S_{\text{lethal}}$ then
7:   $F(G_{j})\leftarrow\text{Lethal}$
8:  else if $S(G_{j})\in S_{\text{free}}$ then
9:   $F(G_{j})\leftarrow\text{Free}$
10:  else
11:   Compute geometric features:
12:   $h\leftarrow\max(|G_{j}^{z}-G_{k}^{z}|)$ {Step height}
13:   $s\leftarrow\arccos(\mathbf{n}\cdot\mathbf{Z})$ {Slope}
14:   $u\leftarrow\log(\text{MSE}(\text{plane},z_{\text{actual}}))$ {Unevenness}
15:   Evaluate GPAM:
16:   $\Psi_{\text{ver}}\leftarrow\llbracket h\leq h_{\max}(V)\rrbracket$
17:   $\Psi_{\text{tre}}\leftarrow\llbracket l_{d}\leq l_{d_{\max}}(V)\rrbracket$
18:   $\Psi_{\text{ove}}\leftarrow\llbracket h_{\text{obj}}>h_{\text{pc}}+h_{\text{Lid}}\rrbracket$
19:   $\Psi_{\text{slo}}\leftarrow\llbracket s\leq\emptyset_{m}(V)\rrbracket$
20:   $\text{GPAM}\leftarrow\llbracket\Psi_{\text{ver}}\land\Psi_{\text{tre}}\land\Psi_{\text{ove}}\land\Psi_{\text{slo}}\rrbracket$
21:   if $\text{GPAM}=0$ then
22:    $F(G_{j})\leftarrow\text{Lethal}$
23:   else
24:    Compute geometric risk:
25:    $\phi_{\text{geo}}\leftarrow\frac{h}{h_{\max}}+\frac{s}{s_{\max}}+\frac{u}{u_{\text{critic}}}$
26:    Assign cost label:
27:    if $\phi_{\text{geo}}\geq\tau_{\text{medium}}(V)$ then
28:     $F(G_{j})\leftarrow\text{M-cost}$
29:    else if $\phi_{\text{geo}}\geq\tau_{\text{low}}(V)$ then
30:     $F(G_{j})\leftarrow\text{L-cost}$
31:    else
32:     $F(G_{j})\leftarrow\text{Free}$
33:    end if
34:   end if
35:  end if
36:end for